OpenStudy (uri):

Stats help: Q.The mean of 12 observations is 3704.One more observation is added and the mean becomes 37.5.Find the value of new observation.

OpenStudy (anonymous):

You can do it!

OpenStudy (anonymous):

Find the total of the operations when there were 12. Then find the total when there were 13.

OpenStudy (anonymous):

What is the difference?

OpenStudy (anonymous):

Since \[ Average = \frac{Total}{Observatioins} \]Then \[ Total = Average \times Observations \]

OpenStudy (anonymous):

find the sum \[\sum x_i= \bar{x}n\]

OpenStudy (anonymous):

\[MEAN x=\frac{ \sum_{}^{}x }{ n }\] total of 12 observations=37.06*12 total of 13 observations=37.5*13 13th observation=? (subtract and find)

OpenStudy (anonymous):

lets call the new observation x \[\bar{x}_2=\frac{\bar{x}_1+x}{n+1}\\\bar{x_1}=37.06,\bar{x}_2=37.5,n+1=13\]

OpenStudy (uri):

I solved the whole question with LatEx and it crashed....

OpenStudy (anonymous):

i mean\[\bar{x}_2=\frac{\bar{x}_1n+x}{n+1}\] \[x=n(\bar{x}_2-\bar{x}_1 )+\bar{x}_2\]

OpenStudy (uri):

I know..I was doing it myself but the page crashed..

OpenStudy (uri):

I so pissed off right now..

OpenStudy (anonymous):

so you cant see anything in LaTex here

OpenStudy (uri):

I know how to do this..i was typing it if i was write.I was at the end of it and it crashed....

OpenStudy (anonymous):

ohhhhh bad...okay ...it does that sometimes ,i guess it has to be reported to the OS feedback

OpenStudy (uri):

Let me do again..oh well.

OpenStudy (uri):

Formula is \[\bar{x} =\frac{ \Sigma x }{ n } \] where n=15 and \[\bar{x} =3704\]

OpenStudy (uri):

So,plugging in the formula \[3704=\frac{ \Sigma x }{ 12 }\]

OpenStudy (uri):

I mistyped above n=12.

OpenStudy (uri):

Then, \[\Sigma x =4448\] let the value of one observation added be *y* Using the same formula again \[37.5=\frac{ 4448+y }{ 13 }\]

OpenStudy (uri):

\[487.5=4448+y\] \[y=4448-487.5\] y=3960.

OpenStudy (uri):

3960.5*

OpenStudy (uri):

@wio Help me c:

OpenStudy (anonymous):

What did you get for your totals?

OpenStudy (uri):

What do you mean?

OpenStudy (anonymous):

is it 37.04 or3704

OpenStudy (uri):

3704.

OpenStudy (anonymous):

\[ 12\times 3706 = ?\\ 13\times 37.5 = ? \]

OpenStudy (uri):

@wio it's 3704* and 12 x 3704=4448 and 13 x 37.5 =487.5

OpenStudy (anonymous):

didnt you mean 37.06 or 3706

OpenStudy (anonymous):

as you have calculated it should b e y=487.5-4448 which is negative

OpenStudy (uri):

Omg sorry i typed wrong,Let me edit the question..Sorry :3

OpenStudy (uri):

@surjithayer we subract the small number with the greater one...No?

OpenStudy (nincompoop):

kick your computer

OpenStudy (uri):

@nincompoop not my computer's fault something is wrong with OS.Happend to many users.

OpenStudy (anonymous):

i think total for 13 observation should be greater than total for 12 observations

OpenStudy (nincompoop):

start doing paid subscription :)

OpenStudy (uri):

I don't think so @surjithayer because the other question i did at my class with my teacher the total for 16 observation was quiet less than the observation for 15..

OpenStudy (nincompoop):

uri, this is the same kind of help you received earlier.

OpenStudy (uri):

your point?

OpenStudy (nincompoop):

you should be able to do this one if you understood the other one :P

OpenStudy (uri):

So,didn't i just do it? I'm asking if i'm right..problem?

OpenStudy (nincompoop):

you had 6 earlier, now you have 12 then 1 is added on both that gave a new mean

OpenStudy (anonymous):

o.k if total for 13 observationis less then 13th observation should be negative

OpenStudy (nincompoop):

(444.48+x)/13=37.5 here's how it is broken down without the fancy statistical jargon 37.04 x 12 = 444.48 multiply it back so you get the summation prior to the unknown. since there were 12 observations, you multiply your initial mean by 12. now we need to find out the unknown (that is +1 observation) in order to get the mean of 37.5. so that is 444. 48 + x then divide it by the final number of observations, 13 (444.48 + x)/13 = 37.5

OpenStudy (uri):

Nin it is 3704 not 37.04..

OpenStudy (nincompoop):

I think that's a typo LOL but if it is really 3,704 then the same kind of process is required for you to find the unknown observation

OpenStudy (uri):

Let me ask a friend

OpenStudy (nincompoop):

okay. it is weird that the new mean is wayyyyyyyyyyyy too low than the initial one. from thousands to tens

OpenStudy (uri):

LOL NIN IT'S 37.4 NOT 3704.My sir doesn't know how to make a decimal point.He wasted my 1 hour..

OpenStudy (nincompoop):

lol told you so… my training in chemistry paid off

OpenStudy (uri):

Omg now i know how to do it..xD

OpenStudy (anonymous):

then it is 37.5*13-37.4*12

OpenStudy (uri):

Let me solve and show. :P

OpenStudy (nincompoop):

yeah so I put 37.04, you just have to make the correction to 37.4

OpenStudy (nincompoop):

you're on the right track

OpenStudy (uri):

Thanks :D

OpenStudy (nincompoop):

always ask your professor what he has written down so you're not wasting a lot of time calculating the wrong information.

OpenStudy (nincompoop):

everyone makes mistakes

OpenStudy (uri):

I do that.I joined late and my class fellows have done this already i just took these photocopied questions from a classmate and did few at class..so didn't know that my teacher makes such a weird typo like WEIRD..0 except a decimal point.

OpenStudy (uri):

@nincompoop look xD

OpenStudy (nincompoop):

lol now you know that small circle is period

OpenStudy (uri):

Lol yess

OpenStudy (anonymous):

there is small circle on is also .you can corelate

OpenStudy (anonymous):

any way you know the procedure to solve such problems.

OpenStudy (uri):

:)